solve using substitution
2y=x+4
5+y=2x

Answer by TimothyLamb(4379)

2y = x + 4
5 + y = 2x
---
put the system of linear equations into standard form
---
x - 2y = -4
2x - y = 5
---
copy and paste the above standard form linear equations in to this solver:
http://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 4.6666667
y = 4.3333333
---
now, solve by substitution:
---
x = 2y - 4
y = 2x - 5
---
x = 2(2x - 5) - 4
x = 4x - 10 - 4
3x = 14
x = 14/3
x = 4.6666666666
y = 2x - 5
y = 2*(4.6666666666) - 5
y = 4.3333333333
---
solution by substitution:
x = 4.6666667
y = 4.3333333
---
Free algebra tutoring live chat:
http://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
http://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
http://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
http://sooeet.com/math/system-of-linear-equations-solver.php
---